OpenStudy (anonymous):

Write an equation for the line that passes through (0,5) and (1,3).

OpenStudy (anonymous):

(y2-y1)/(x2-x1) (3-5)/(1-0)=-2/1 =-2x

OpenStudy (anonymous):

5=-2(0)+b 5=0+b 5=b

OpenStudy (anonymous):

y=-2x+5

OpenStudy (anonymous):

does that help?